# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 6. Pure infiniteness
Let $y\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\setminus\{0\}$. We show that $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})y$ contains an infinite idempotent. Let $0<\varepsilon<\frac{1}{192}$. Arguing as in the beginning of proof of Theorem 6.7 we may assume that $\|\operatorname{\mathbb{E}}^{\mathcal{R}}_{X}(y)\|=1$ and that there is $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with $\|y-i_{\mathcal{R}}(f)\|_{\mathcal{R}}<\varepsilon$ and $\|E^{\mathcal{R}}_{X}(i_{\mathcal{R}}(f))\|=\|q_{X}(f|_{X})\|=1$. By Lemma 5.8 there is a contractive $b\in C_{c}(X)$ and a bump function $a\in C_{c}(X)$ such that $bf|_{X}\in C_{c}(X)^{+}$ and
@@
This implies that $\|i_{\mathcal{R}}(p)-i_{\mathcal{R}}(u^{*})(by)i_{\mathcal{R}}(u)\|_{\mathcal{R}}<1$ and therefore $i_{\mathcal{R}}(u^{*})(by)i_{\mathcal{R}}(u)$ is invertible in the Banach algebra $i_{\mathcal{R}}(p)F_{\mathcal{R}}(\mathcal{G},\mathcal{L})i_{\mathcal{R}}(p)$. We denote by $c$ its inverse and we put $w:=c\cdot i_{\mathcal{R}}(u^{*})(by)\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})y$. Then $wi_{\mathcal{R}}(u)=i_{\mathcal{R}}(p)$ is an infinite idempotent and therefore the equivalent idempotent $i_{\mathcal{R}}(u)w\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})y$ is also infinite. ∎
@@
Let $(\mathcal{G},\mathcal{L})$ be a twisted étale groupoid where $\mathcal{G}$ is topologically free and minimal. If $\mathcal{G}$ is locally contracting with respect to a unital subsemigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$, then every every $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ of $(\mathcal{G},\mathcal{L})$ is purely infinite simple.
@@ 7.1. Crossed products by twisted partial group actions
We use the transformation groupoid model from [BKM23] to apply our results to twisted partial group actions. In particular, we will generalize and extend some of the results of [BK24] from global to partial actions.
We fix a (topological) partial action of a discrete group $G$ on a locally compact Hausdorff space $X$. This is a map $\theta:G\to\operatorname{PHomeo}(X)$ such that $\theta_{1}=\operatorname{id}_{X}$, $\theta_{t}^{-1}=\theta_{t^{-1}}$ and the partial homeomorphism $\theta_{ts}$ extends $\theta_{t}\circ\theta_{s}$ for all $t,s\in G$. Thus for each $t\in G$ we have a homeomorphism $\theta_{t}:X_{t}\to X_{t^{-1}}$ between two open subsets of $X$ and for each $s\in G$ the natural domain $\theta_{t^{-1}}(X_{s^{-1}}\cap X_{t})=X_{t^{-1}}\cap X_{(st)^{-1}}$ for $\theta_{s}\circ\theta_{t}$ is contained in $X_{(st)^{-1}}$. Fixing $\theta$ is equivalent to fixing an algebraic partial action $\alpha:G\to\operatorname{PAut}(C_{0}(X))$ where $\alpha_{t}(a):=a\circ\theta_{t}^{-1}$, $a\in C_{0}(X_{t^{-1}})$, $t\in S$. We refer to $\theta$ as dual to $\alpha$.
We also fix a (normalized) 2-cocycle of the partial action $\alpha$ (or of $\theta$), which is a family of maps $u(s,t):G\times G\to C_{u}(X_{s}\cap X_{st})$, $s,t\in G$, satisfying $u(1,t)=u(t,1)=1$ for $t\in G$, and
$\alpha_{r}\big{(}au(s,t)\big{)}u(r,st)=\alpha_{r}(a)u(r,s)u(rs,t),\qquad a\in C_{0}(X_{r^{-1}}\cap X_{s}\cap X_{st}),\ r,s,t\in G.$
Definition 7.1 (cf. [BKM23, Definition 7.3] and [BK24, Definitions 2.1, 4.12]).
@@
with operations
@@
for $f,g\in\ell^{1}(\alpha,u)$ and $s,t,r\in G$. For any class $\mathcal{R}$ of representations of $F(\alpha,u)$ we denote by $F_{\mathcal{R}}(\alpha,u)$ the Hausdorff completion of $F(\alpha,u)$ in the seminorm $\|f\|_{\mathcal{R}}:=\sup\{\|\psi(f)\|:\psi\in\mathcal{R}\}$, $f\in F(\alpha,u)$. For any non-empty $P\subseteq[1,\infty]$ we define the $L^{P}$-crossed product for $(\theta,u)$ as the completion $F^{P}(\alpha,u):=\overline{F(\alpha,u)}^{\|\cdot\|_{L_{P}}}$, where $\|f\|_{L^{P}}$ is the supremum of $\|\psi(f)\|$ over representations $\psi:F(\alpha,u)\to B(L^{p}(\mu))$ where $p\in P$ and $\mu$ is a measure, and if $\mathbb{F}=\mathbb{R}$ we also assume that $\psi|_{C_{0}(X)}$ acts by multiplication operators.
Remark 7.2.
By [BKM23, Theorem 7.9] representations of $F(\alpha,u)$ can be described in terms of covariant representations of $(\alpha,u)$, and when $\mathbb{F}=\mathbb{C}$, then $F^{P}(\theta,u)$ is universal for all representations of $F(\alpha,u)$ in all $L^{p}$-operator algebras for $p\in P$. Moreover we have a natural groupoid model for the algebras $F_{\mathcal{R}}(\alpha,u)$. Namely, the transformation groupoid for $\theta$ is
@@
The twist associated to $u$ is given by the trivial line bundle $\mathcal{L}_{u}=\mathcal{G}_{\theta}\times\mathbb{F}$ and the $2$-cocycle $\sigma_{u}$ for $\mathcal{G}_{\theta}$ given by $\sigma_{u}(s,\theta_{t}(x)),t,x)=u(s,t)(\theta_{st}(x))$. Let $S_{\theta}$ be the inverse semigroup consisting of bisections $\{t\}\times U\subseteq\mathcal{G}_{\theta}$ where $U\subseteq X_{t^{-1}}$ is open. Then by [BKM23, Theorem 7.9] the map $C_{c}(\mathcal{G}_{\theta},\mathcal{L}_{u})\ni\hat{f}\mapsto f\in C_{c}(G,C_{c}(X))$ where $f(t)(x):=\hat{f}(t,\theta_{t^{-1}}(x))$ extends to an isometric isomorphism
@@
Thus for any class $\mathcal{R}$ of representations of $F(\alpha,u)$ there is a class $\mathcal{R}_{0}$ of representations of $F^{S_{\theta}}(\mathcal{G}_{\theta},\mathcal{L}_{u})$ such that (7.13) induces an isometric isomorphism $F^{S_{\theta}}_{\mathcal{R}_{0}}(\mathcal{G}_{\theta},\mathcal{L}_{u})\cong F_{\mathcal{R}}(\alpha,u)$
@@
For any Banach algebra of the form $F_{\mathcal{R}}(\alpha,u)$ the following are equivalent:
@@
Using the groupoid model from Remark 7.2 the assertion follows from Proposition 3.6, see also Remark 3.8. ∎
@@
If the equivalent conditions in Lemma 7.3 hold then we call $F_{\mathcal{R}}(\alpha,u)$ a Banach algebra crossed product for $(\alpha,u)$. If in addition $\ker j_{\mathcal{R}}=\bigcap_{t\in G}\ker E_{t}^{\mathcal{R}}=\{0\}$, we say that $F_{\mathcal{R}}(\alpha,u)$ is a reduced Banach algebra crossed product.
For each $p\in[1,+\infty]$ the representation (2.3) composed with the isomorphism (7.13) gives a representation $\Lambda_{p}:F(\alpha,u)\to B(\ell^{p}(\mathcal{G}_{\theta}))$, where
@@
For any non-empty $P\subseteq[1,+\infty]$ we call $F^{P}_{\operatorname{r}}(\alpha,u):=F_{\{\Lambda_{p}:p\in P\}}(\alpha,u)$ the reduced $L^{P}$-crossed product. We define the full $L^{P}$-crossed product $F^{P}(\alpha,u)$ as $F_{\mathcal{R}}(\alpha,u)$ where $\mathcal{R}$ is the class of all representations $\psi$ of $F(\alpha,u)$ on $L^{p}$-spaces $L^{p}(\mu)$, for $p\in P$, and if $\mathbb{F}=\mathbb{R}$ we additionally assume that $\psi(C_{0}(X))$ consists of operators of multiplication. We denote by $\Lambda_{P}:F^{P}(\alpha,u)\to F^{P}_{\operatorname{r}}(\alpha,u)$ the canonical representation.
By (7.13) and Proposition 3.15, the algebras $F^{P}_{\operatorname{r}}(\alpha,u)$ are reduced crossed product isometrically isomorphic to $F^{P}_{\operatorname{r}}(\mathcal{G}_{\theta},\mathcal{L}_{u})$. The above definitions agree with those introduced in [BK24] for full actions and $\mathbb{F}=\mathbb{C}$.
@@
The partial action $\theta$ is topologically free if $\{x\in X_{t^{-1}}:\theta_{t}(x)=x\}$ has empty interior for every $t\in G\setminus\{1\}$. A set $U\subseteq X$ is $\theta$-invariant if $h_{t}(U\cap X_{t^{-1}})\subseteq U$ for all $t\in G$. The partial action $\theta$ is minimal if there are no non-trivial $\theta$-invariant open sets.
@@
It is obvious that the partial action $\theta$ is minimal if and only if the associated transformation groupoid $\mathcal{G}_{\theta}$ is minimal. It is also easy to see from Remark 5.6 that the partial action $\theta$ is topologically free if and only if $\mathcal{G}_{\theta}$ is topologically free.
When the twist $u$ is trivial, we omit writing it, and when $H\subseteq G$ is a subgroup then $(\alpha|_{H},u|_{H})$ denotes the obvious restriction of the twisted action $(\alpha,u)$ of $G$ to a twisted action of $H$.
@@
Let $(\alpha,u)$ be a twisted partial action $(\alpha,u)$ of a discrete group $G$ on the algebra $C_{0}(X)$, with dual partial action $\theta$ of $G$ on $X$. Let $P\subseteq[1,\infty]$ be a non-empty set and let $F_{\mathcal{R}}(\alpha,u)$ be any reduced crossed product Banach algebra. The following are equivalent:
@@
If the above equivalent conditions hold then $C_{0}(X)$ detects ideals in $F_{\mathcal{R}}(\alpha,u)$ and so $F_{\mathcal{R}}(\alpha,u)$ is simple if and only if $\theta$ is minimal.
@@
Extending isomorphism (7.13) we may view a reduced crossed product Banach algebra $F_{\mathcal{R}}(\alpha,u)$ of $(\alpha,u)$ as a reduced groupoid Banach algebra of $(\mathcal{G}_{\theta},\mathcal{L}_{u})$. Since $\mathcal{G}_{\theta}$ is Hausdorff, and topological freeness and minimality of $\theta$ are equivalent to topological freeness and minimality of $\mathcal{G}_{\theta}$ the assertion follows from Proposition 5.11 and Theorems 5.10, 5.13, 5.16, cf. also Theorem A and Corollary B. Formally to get from (4) to (1) one needs to modify the proof of Theorem 5.16 by replacing $U$ with its intersection with some $\{t\}\times X_{t^{-1}}$, $t\in G$. ∎
Moving towards pure infiniteness we generalize to the partial action case the definitions of $n$-filling and local boundary actions introduced in [JR00] and [LS96] respectively.
@@
For $n\geq 2$ we say that the partial action $\theta$ is $n$-filling if $X$ is compact, infinite and for any non-empty open set $U\subseteq X$ there are $t_{1},\ldots,t_{n}\in G$ with $\bigcup_{i=1}^{n}\theta_{t_{i}}(X_{t_{i}^{-1}}\cap U_{i})=X$.
@@
For full actions, $2$-filling is equivalent to being a strong boundary action introduced by Laca and Spielberg [LS96]. In general, the $n$-filling condition for the action $\theta$ is stronger than the $n$-filling condition for the groupoid $\mathcal{G}_{\theta}$, see [Suz17].
@@
If the action $\theta$ is locally boundary then the groupoid $\mathcal{G}_{\theta}$ is locally contracting (in fact the latter notion was inspired by the former, see [A-D97]).
@@
Let $(\alpha,u)$ be a twisted partial action of a discrete group $G$ on $C_{0}(X)$. If the dual action $\theta$ is topologically free, minimal and either $n$-filling or locally boundary, then every reduced twisted crossed product Banach algebra $F_{\mathcal{R}}(\alpha,u)$ is purely infinite simple.
@@
Extending isomorphism (7.13) we may view $F_{\mathcal{R}}(\alpha,u)$ as a reduced groupoid Banach algebra of $(\mathcal{G}_{\theta},\mathcal{L}_{u})$. Since $\mathcal{G}_{\theta}$ is Hausdorff this is also an essential groupoid Banach algebra, and $\mathcal{G}_{\theta}$ is topologically free and minimal. Clearly $n$-filling, with $n\in\mathbb{N}$, implies $n$-filling of $\mathcal{G}_{\theta}$ with respect to $S_{\theta}$, and if $\theta$ is a local boundary action then $\mathcal{G}_{\theta}$ is locally contracting with respect to $S_{\theta}$. The assertion therefore follows from Theorem 6.7 and Theorem 6.13. ∎
@@
The above theorem applies to a number of natural group actions on boundaries of various type, cf. [A-D97] and [JR00], to actions modeling Ruelle algebras arising from certain Smale spaces, see [LS96, Examples 14.1, 14.2], and to many actions of any countable non-amenable exact group on a product of any connected closed topological manifold with a Cantor set, see [Suz17, Theorem 4.1]. If $X$ is totally disconnected and every compact open set is $\theta$-paradoxical, in the sense of [GS14, Definition 4.3], then $\theta$ is a local boundary action. Thus Theorem 7.13 applies to partial actions described in [GS14, Section 5] such as those modeling Cuntz algebras, simple Cuntz–Krieger algebras and algebras of integral domains.

### v3

@@ 6. Pure infiniteness
Let $y\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\setminus\{0\}$. We show that $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})y$ contains an infinite idempotent. Let $0<\varepsilon<\frac{1}{192}$. Arguing as in the beginning of the proof of Theorem 6.7, we may assume that $\|\operatorname{\mathbb{E}}^{\mathcal{R}}_{X}(y)\|=1$ and that there is $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with $\|y-i_{\mathcal{R}}(f)\|_{\mathcal{R}}<\varepsilon$ and $\|E^{\mathcal{R}}_{X}(i_{\mathcal{R}}(f))\|=\|q_{X}(f|_{X})\|=1$. By Lemma 5.8 there is a contractive $b\in C_{c}(X)$ and a bump function $a\in C_{c}(X)$ such that $bf|_{X}\in C_{c}(X)^{+}$ and
@@
This implies that $\|i_{\mathcal{R}}(p)-i_{\mathcal{R}}(u^{*})(by)i_{\mathcal{R}}(u)\|_{\mathcal{R}}<1$ and therefore $i_{\mathcal{R}}(u^{*})(by)i_{\mathcal{R}}(u)$ is invertible in the Banach algebra $i_{\mathcal{R}}(p)F_{\mathcal{R}}(\mathcal{G},\mathcal{L})i_{\mathcal{R}}(p)$. We denote its inverse by $c$ and we put $w:=c\cdot i_{\mathcal{R}}(u^{*})(by)\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})y$. Then $wi_{\mathcal{R}}(u)=i_{\mathcal{R}}(p)$ is an infinite idempotent and therefore the equivalent idempotent $i_{\mathcal{R}}(u)w\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})y$ is also infinite. ∎
@@
Let $(\mathcal{G},\mathcal{L})$ be a twisted étale groupoid where $\mathcal{G}$ is topologically free and minimal. If $\mathcal{G}$ is locally contracting with respect to a unital subsemigroup $S\subseteq S(\mathcal{L})$ covering $\mathcal{G}$, then every $S$-graded essential Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ of $(\mathcal{G},\mathcal{L})$ is purely infinite simple.
@@ 7.1. Crossed products by twisted partial group actions
We use a transformation groupoid model to apply our results to twisted partial group actions. In particular, we will generalize and extend some of the results of [BK24] from global to partial actions.
We fix a (topological) partial action of a discrete group $G$ on a locally compact Hausdorff space $X$. This is a map $\theta:G\to\operatorname{PHomeo}(X)$ such that $\theta_{1}=\operatorname{id}_{X}$, $\theta_{t}^{-1}=\theta_{t^{-1}}$ and the partial homeomorphism $\theta_{ts}$ extends $\theta_{t}\circ\theta_{s}$ for all $t,s\in G$. Thus for each $t\in G$ we have a homeomorphism $\theta_{t}:X_{t^{-1}}\to X_{t}$ between two open subsets of $X$ and for each $s\in G$ the natural domain $\theta_{t^{-1}}(X_{s^{-1}}\cap X_{t})=X_{t^{-1}}\cap X_{(st)^{-1}}$ for $\theta_{s}\circ\theta_{t}$ is contained in $X_{(st)^{-1}}$. Fixing $\theta$ is equivalent to fixing an algebraic partial action $\alpha:G\to\operatorname{PAut}(C_{0}(X))$ where $\alpha_{t}(a):=a\circ\theta_{t}^{-1}$, $a\in C_{0}(X_{t^{-1}})$, $t\in S$. We refer to $\theta$ as dual to $\alpha$.
We also fix a 2-cocycle of the partial action $\alpha$ (or of $\theta$), which is a family of maps $u(s,t):G\times G\to C_{u}(X_{s}\cap X_{st})$, $s,t\in G$, satisfying $u(1,t)=u(t,1)=1$ for $t\in G$, and
$\alpha_{r}\big(au(s,t)\big)u(r,st)=\alpha_{r}(a)u(r,s)u(rs,t),\qquad a\in C_{0}(X_{r^{-1}}\cap X_{s}\cap X_{st}),\ r,s,t\in G,$
see [Exe97].
Definition 7.1 (cf. [BK24, Definitions 2.1, 4.12]).
@@
with $\ell^{1}$-norm from $\ell^{1}(G,C_{0}(X))$ and operations
@@
for $f,g\in F(\alpha,u)$ and $s,t,r\in G$. For any class $\mathcal{R}$ of representations of $F(\alpha,u)$, we denote the Hausdorff completion of $F(\alpha,u)$ in the seminorm $\|f\|_{\mathcal{R}}:=\sup\{\|\psi(f)\|:\psi\in\mathcal{R}\}$, $f\in F(\alpha,u)$, by $F_{\mathcal{R}}(\alpha,u)$. For any non-empty $P\subseteq[1,\infty]$ we define the $L^{P}$-crossed product for $(\theta,u)$ as the completion $F^{P}(\alpha,u):=\overline{F(\alpha,u)}^{\|\cdot\|_{L_{P}}}$, where $\|f\|_{L^{P}}$ is the supremum of $\|\psi(f)\|$ over representations $\psi:F(\alpha,u)\to B(L^{p}(\mu))$, where $p\in P$ and $\mu$ is a measure, and if $\mathbb{F}=\mathbb{R}$ we also assume that $\psi|_{C_{c}(X)}$ is positive.
We briefly describe how to view the algebras $F_{\mathcal{R}}(\alpha,u)$ as groupoid Banach algebras. More details, as well as description of representations of $F(\alpha,u)$ in terms of covariant representations of $(\alpha,u)$, can be found in an earlier preprint version of [BKM25], preserved as arXiv:2303.09997v3, see subsection 7.1 therein.
Example 7.2.
The transformation groupoid for $\theta$ is
@@
The twist associated to $u$ is given by the trivial line bundle $\mathcal{L}_{u}=\mathcal{G}_{\theta}\times\mathbb{F}$ and the $2$-cocycle $\sigma_{u}$ for $\mathcal{G}_{\theta}$ given by
$\sigma_{u}((s,\theta_{t}(x)),(t,x))=u(s,t)(\theta_{st}(x)).$
Let $S_{\theta}$ be the inverse semigroup consisting of bisections $\{t\}\times U\subseteq\mathcal{G}_{\theta}$ where $U\subseteq X_{t^{-1}}$ is open. Then one sees that the map $C_{c}(\mathcal{G}_{\theta},\mathcal{L}_{u})\ni\hat{f}\mapsto f\in C_{c}(G,C_{c}(X))$ where $f(t)(x):=\hat{f}(t,\theta_{t^{-1}}(x))$ extends to an isometric isomorphism
@@
Thus for any class $\mathcal{R}$ of representations of $F(\alpha,u)$ there is a class $\mathcal{R}_{0}$ of representations of $F^{S_{\theta}}(\mathcal{G}_{\theta},\mathcal{L}_{u})$ such that (7.12) induces an isometric isomorphism $F^{S_{\theta}}_{\mathcal{R}_{0}}(\mathcal{G}_{\theta},\mathcal{L}_{u})\cong F_{\mathcal{R}}(\alpha,u)$.
@@
For any Banach algebra of the form $F_{\mathcal{R}}(\alpha,u)$, the following are equivalent:
@@
Using the groupoid model from Example 7.2, the assertion follows from Proposition 3.6, see also Remark 3.8. ∎
@@
If the equivalent conditions in Lemma 7.3 hold, then we call $F_{\mathcal{R}}(\alpha,u)$ a Banach algebra crossed product for $(\alpha,u)$. If in addition $\ker j_{\mathcal{R}}=\bigcap_{t\in G}\ker E_{t}^{\mathcal{R}}=\{0\}$, we say that $F_{\mathcal{R}}(\alpha,u)$ is a reduced Banach algebra crossed product.
For each $p\in[1,+\infty]$, the representation (2.2) composed with the isomorphism (7.12) gives a representation $\Lambda_{p}:F(\alpha,u)\to B(\ell^{p}(\mathcal{G}_{\theta}))$, where
@@
For any non-empty $P\subseteq[1,+\infty]$, we call $F^{P}_{\operatorname{r}}(\alpha,u):=F_{\{\Lambda_{p}:p\in P\}}(\alpha,u)$ the reduced $L^{P}$-crossed product. We define the full $L^{P}$-crossed product $F^{P}(\alpha,u)$ as $F_{\mathcal{R}}(\alpha,u)$ where $\mathcal{R}$ is the class of all representations $\psi$ of $F(\alpha,u)$ on $L^{p}$-spaces $L^{p}(\mu)$, for $p\in P$, and if $\mathbb{F}=\mathbb{R}$ we additionally assume that $\psi(C_{0}(X))$ consists of multiplication operators. We let $\Lambda_{P}:F^{P}(\alpha,u)\to F^{P}_{\operatorname{r}}(\alpha,u)$ be the canonical representation.
By (7.12) and Proposition 3.15, the algebras $F^{P}_{\operatorname{r}}(\alpha,u)$ are reduced crossed products isometrically isomorphic to $F^{P}_{\operatorname{r}}(\mathcal{G}_{\theta},\mathcal{L}_{u})$. The above definitions agree with those introduced in [BK24] for full actions and $\mathbb{F}=\mathbb{C}$.
@@
A partial action $\theta$ is topologically free if $\{x\in X_{t^{-1}}:\theta_{t}(x)=x\}$ has empty interior for every $t\in G\setminus\{1\}$. A set $U\subseteq X$ is $\theta$-invariant if $h_{t}(U\cap X_{t^{-1}})\subseteq U$ for all $t\in G$. A partial action $\theta$ is minimal if there are no non-trivial $\theta$-invariant open sets.
@@
It is obvious that a partial action $\theta$ is minimal if and only if the associated transformation groupoid $\mathcal{G}_{\theta}$ is minimal. It is also easy to see from Remark 5.6 that the partial action $\theta$ is topologically free if and only if $\mathcal{G}_{\theta}$ is topologically free.
If the twist $u$ is trivial, then we omit writing it. For a subgroup $H\subseteq G$ we denote the obvious restriction of the twisted action $(\alpha,u)$ of $G$ to a twisted action of $H$ by $(\alpha|_{H},u|_{H})$.
@@
Let $(\alpha,u)$ be a twisted partial action of a discrete group $G$ on the algebra $C_{0}(X)$, with dual partial action $\theta$ of $G$ on $X$. Let $P\subseteq[1,\infty]$ be a non-empty set and let $F_{\mathcal{R}}(\alpha,u)$ be any reduced crossed product Banach algebra. The following are equivalent:
@@
If the above equivalent conditions hold, then $C_{0}(X)$ detects ideals in $F_{\mathcal{R}}(\alpha,u)$ and so $F_{\mathcal{R}}(\alpha,u)$ is simple if and only if $\theta$ is minimal.
@@
Extending isomorphism (7.12), we may view a reduced crossed product Banach algebra $F_{\mathcal{R}}(\alpha,u)$ of $(\alpha,u)$ as an $S_{\theta}$-graded reduced groupoid Banach algebra of $(\mathcal{G}_{\theta},\mathcal{L}_{u})$. As noted in Remark 7.7, topological freeness and minimality of $\theta$ are equivalent to topological freeness and minimality of $\mathcal{G}_{\theta}$. The groupoid $\mathcal{G}_{\theta}$ is Hausdorff. Hence the assertion follows from Proposition 5.11 and Theorems 5.10, 5.13, 5.16, cf. also Theorem A and Corollary B. Formally, to get from (4) to (1), one needs to modify the proof of Theorem 5.16 by replacing $U$ by its intersection with some $\{t\}\times X_{t^{-1}}$, $t\in G$. ∎
Moving towards pure infiniteness, we generalize the definitions of $n$-filling and local boundary actions, introduced in [JR00] and [LS96] respectively, to the partial action case.
@@
For $n\geq 2$, we say that a partial action $\theta$ is $n$-filling if $X$ is compact, infinite and for any non-empty open set $U\subseteq X$ there are $t_{1},\ldots,t_{n}\in G$ with $\bigcup_{i=1}^{n}\theta_{t_{i}}(X_{t_{i}^{-1}}\cap U_{i})=X$.
@@
For full actions, 22-filling is equivalent to being a strong boundary action introduced by Laca and Spielberg [LS96]. In general, the $n$-filling condition for the action $\theta$ is stronger than the $n$-filling condition for the groupoid $\mathcal{G}_{\theta}$, see [Suz17].
@@
If $\theta$ is a local boundary action, then the groupoid $\mathcal{G}_{\theta}$ is locally contracting (in fact the latter notion was inspired by the former, see [A-D97]).
@@
Let $(\alpha,u)$ be a twisted partial action of a discrete group $G$ on $C_{0}(X)$. If the dual action $\theta$ is topologically free, minimal and either $n$-filling or a local boundary action, then every reduced twisted crossed product Banach algebra $F_{\mathcal{R}}(\alpha,u)$ is purely infinite simple.
@@
Extending isomorphism (7.12), we may view $F_{\mathcal{R}}(\alpha,u)$ as a reduced groupoid Banach algebra of $(\mathcal{G}_{\theta},\mathcal{L}_{u})$. Since $\mathcal{G}_{\theta}$ is Hausdorff, this is also an essential groupoid Banach algebra, and $\mathcal{G}_{\theta}$ is topologically free and minimal. Clearly, if $\theta$ is $n$-filling then $\mathcal{G}_{\theta}$ is $n$-filling with respect to $S_{\theta}$, and if $\theta$ is a local boundary action then $\mathcal{G}_{\theta}$ is locally contracting with respect to $S_{\theta}$. The assertion therefore follows from Theorem 6.7 and Theorem 6.13. ∎
@@
Theorem 7.13 applies to a number of natural group actions on boundaries of various type, cf. [A-D97] and [JR00], to actions modeling Ruelle algebras arising from certain Smale spaces, see [LS96, Examples 14.1, 14.2], and to many actions of any countable non-amenable exact group on a product of any connected closed topological manifold with a Cantor set, see [Suz17, Theorem 4.1]. If $X$ is totally disconnected and every compact open set is $\theta$-paradoxical, in the sense of [GS14, Definition 4.3], then $\theta$ is a local boundary action. Thus Theorem 7.13 applies to partial actions described in [GS14, Section 5] such as those modeling Cuntz algebras, simple Cuntz–Krieger algebras and algebras of integral domains.